OpenStudy (anonymous):

help Equation M: 3y = 3x + 6 Equation P: y = x + 2 Which of the following best describes the solution to the given set of equations? A. No solution B. One solution C. Two solutions D. Infinite solutions

OpenStudy (kidrah69):

here I used substitution now just solve the problem 3(x+2)= 3x + 6

OpenStudy (anonymous):

hey i need to know how many solutions there is!

OpenStudy (kidrah69):

you will know once you solve the problem

OpenStudy (anonymous):

but i don't now how to solve it!

OpenStudy (kidrah69):

If you get 0=0 it means that there are infinite soltuion so any of the same numbers equalling themself so 3=3 4=4 anything One solution is like x=# no solution is 0=1 so a number wont equal another number. to solve the problem first distribute the 3

OpenStudy (anonymous):

3x and 6

OpenStudy (kidrah69):

3x+6=3x+6 combine like terms

OpenStudy (anonymous):

+6 +6 =9

OpenStudy (anonymous):

opps i ment minuse

OpenStudy (kidrah69):

no we want to isolate x yep so 6-6

OpenStudy (anonymous):

0

OpenStudy (kidrah69):

3x-3x=?

OpenStudy (anonymous):

0x

OpenStudy (anonymous):

0

OpenStudy (kidrah69):

0=0 so how many soltuions are there?

OpenStudy (anonymous):

many

OpenStudy (anonymous):

?

OpenStudy (kidrah69):

yes! :)

OpenStudy (kidrah69):

good job!

OpenStudy (kidrah69):

also welcome to openstudy